### v1

# Neutrino Telescope Event Classification On Quantum Computers

Abstract
[…]
I Introduction
In recent years, the use of quantum computers for classification tasks has garnered significant attention. On the one hand, there has been theoretical exploration into the separation between quantum and classical learners; on the other, practical implementations have aimed to demonstrate that quantum machine learning (QML) models [1] can tackle complex classification problems. Among various QML approaches, quantum kernel methods are particularly prominent, and have been applied to a range of domains including cosmology [2], phase classification in condensed matter systems [3, 4], and satellite image analysis [5]. Experimental realizations of quantum kernel models have also been reported [6, 7]. Similarly, quantum convolutional methods [8], which are motivated by the success of convolutional neural networks, have been used in image processing [9] and recognition [10], and particle physics applications [11].
In high-energy physics, numerous classification tasks have been addressed using QML models [12]. For example, Refs. [13, 14, 15] develop quantum computing algorithms to separate signal, e.g., Higgs events, from background in LHC and, in Ref. [16], the authors have studied the encoding of neutrino telescope data in quantum systems. However, one of the primary challenges facing QML is the encoding of large feature spaces [17], a critical bottleneck in high-energy physics applications, where classification problems often involve vast amounts of information. To mitigate this, various dimensionality reduction techniques such as autoencoders [18], principal component analysis (PCA) [19], and the use of high-level features [20, 21, 14] have been explored.
Nevertheless, when it comes to classifying different neutrino interactions within neutrino telescopes, such as the IceCube Neutrino Observatory in the South Pole [22] or the KM3NeT detector in the Mediterranean [23], the standard approach has relied on graph neural networks (GNNs) that utilize full detector information [24, 25].
For such data-rich events, executing these classification tasks on current quantum hardware is unfeasible due to limited qubit counts and high noise levels, which hinder the practical deployment of quantum GNNs. Finding a suitable preprocessing approach which is still expressive enough to capture the underlying complexity of the full detector data while not exhausting current technological limits is crucial to the success of the QML algorithm. Among the various approaches explored, we highlight below the most successful method.
In this work, we propose a preprocessing strategy for detector data, inspired by the underlying physics and geometry of the problem, that enables state-of-the-art classification performance with a drastically reduced number of features. This reduction makes it feasible to solve the classification task using a small number of qubits, marking, to the best of our knowledge, the first time this problem has been successfully addressed on a quantum computer. We validate our approach not only through simulations but also via implementation on real quantum hardware. The goal of our quantum classifiers is to separate muons versus hadronic and electrotromagnetic showers, providing neutrino telescopes flavor separation. To do so, we will use supervised-learning, where we will simulate muon and electron neutrino event signatures in neutrino telescopes, such as IceCube.
Figure 1: Moment of inertia representations and event histogram binned in energy decade On the left, the two data categories of interest, (a) tracks and (b) cascades, are shown. Schematic representations of the typical shapes are superimposed on each, illustrating the rationale behind employing moments of inertia as a distinguishing feature. In the event displays, each circle represents an Optical Module (OM): the size of the circle indicates the number of photons detected by that OM (larger circles correspond to more photons), and the color shows the photon arrival time (with red for earlier and green for later arrivals). On the right (c), the 2D histogram displays the number of photons per event, grouped by energy bins. For each energy decade, 1000 events were generated — half are tracks and half are cascades.
II Dataset
To demonstrate the application of quantum classification algorithms to neutrino astronomy, we must first simulate the response of a neutrino telescope. For definiteness, we will discuss the case of IceCube, though our approach is readily extendable to other neutrino telescopes geometries and detection media. IceCube detects neutrinos by observing Cherenkov light emitted by charged particles from neutrino interactions in the ice. The Cherenkov photons are recorded by Optical Modules (OMs) — each containing a photomultiplier tube and on-board digitizer [22] — that are arranged in a three dimensional hexagonal lattice in the Antartic ice deployed between 2.5 and 3 kilometers deep. The photomultipler converts incident light into current, which is then read out by an onboard-digitizer, providing us with time-stamped light detection. The simulation of the particle interactions, light propagation, and detector is done using Prometheus [26], an open source neutrino telescope simulation software.
The pattern of observed light typically falls into two morphological categories: tracks and cascades. Tracks are long cylindrically-shaped light depositions created by high-energy muons, which typically travel several kilometers, see Figure 1(a). These muons are predominantly produced in cosmic-ray air showers and by charged-current muon neutrino interactions in the ice, with a sub-dominant contribution from tau neutrinos. Cascades, in contrast, appear as nearly spherical light depositions. They result from charged-current interactions of electron and tau neutrinos, as well as neutral-current interactions of all neutrino flavors. The separation between these two categories can be done in multiple ways, e.g., one could measure the energy loss per unit length [27], transform the photon hits into a 4-dimensional image and use a convolutional neural network [28] or, as recently done, one could embed the photons into a graph [24]. The latter approaches do not take advantage of physics intuition and instead lean on recent machine learning advances, with the hope that they will prove beneficial. Instead, in this work, we use geometrical intuition to separate these two categories by their shapes [29].
As illustrated in Figure 1 (a) and (b), tracks and cascades exhibit distinctly different geometrical signatures. We use the moment of inertia to separate these signatures, as it captures the amount of light observed in addition to the spatial structure. The moment inertia of cascades is spherical-like, while tracks are more akin to cylinders, allowing for their separation through this variable. Later, we will see that we can also add the distance traveled by the center of mass to further differentiate between these morphologies.
For this work, we simulated neutrinos interacting in IceCube in four energy ranges: 100 GeV - 1 TeV, 1 TeV - 10 TeV, 10 TeV - 100 TeV, and 100 TeV - 1 PeV. For each dataset, we produced 500 tracks and 500 cascades, by injecting muon neutrinos and electron neutrinos, respectively.
The number of detected photons per generated track or cascade vary widely and depend on the energy of the incident neutrino; see Figure 1(c). The task of encoding this data into a quantum system is in itself a difficult challenge as there can be up to one million features, if we take each detected photon to be a single feature, exceeding the data encoding capabilities of current quantum computers.
To accommodate current technological constraints, we preprocess the data by leveraging the geometric differences between tracks and cascades following an inductive bias.
Thus, instead of encoding every individual photon or summary statistics as done in Refs. [28, 30, 24, 16, 31], we encode only the shape of the event by computing the moment of inertia tensor for every track and cascade. The moment of inertia tensor is computed as
$I_{ij}=\sum_{k=1}^{N_{\rm OM}}q_{k}\left(\left\|\vec{r}_{k}\right\|^{2}\delta_{ij}-x_{i}^{(k)}x_{j}^{(k)}\right),$ (1)
where $q_{k}$ is the number of photons detected in the $k$-th OM, $\vec{r}_{k}$ is the vector from the OM to the center-of-mass of the track or the cascade, $x_{i}^{(k)}$ is the $i$-th component of $\vec{r}_{k}$, and $N_{\rm OM}$ is the number of OMs that detected light. Additionally, the distance traversed by the center-of-mass as the OMs detect light is also an important differentiator between cascades and tracks. We define the variable CoM as the change in position of the center-of-mass from the moment when 1/4 of the total light was detected to the moment when 3/4 of the light was observed. This choice is motivated by the fact that tracks travel long distances in the detector, while cascades are more localized. To account for variations in event brightness, the CoM distance is normalized by the maximum CoM distance observed in the dataset. Finally, we take the eigenvalues of $I_{ij}$ and assign the smallest to largest eigenvalues to variables denoted by $I_{0}$, $I_{1}$, and $I_{2}$. Since the total number of hits can vary significantly across events, the three eigenvalues are normalized on an event by event basis, as only their relative values carry meaningful geometric information.
Thus our data set reduces every data entry, which potentially contains millions of photons, into four features: CoM, $I_{0}$, $I_{1}$, and $I_{2}$. The distribution of these four features across our data set for the four energy bins is shown in Figure 2. As can be seen in this figure tracks and cascade populations can be separated using this information. For comparison, other preprocessing steps used in neutrino telescope data include computing the quantiles of the light observed in each optical module [28] and associating it to a graph [24]. This approach is numerically more expensive than ours, which scales linearly, $\mathcal{O}(N_{OM})$, with the number of optical modules and is independent of the number of photons per module. In contrast, the quantile calculation scales as $\mathcal{O}(N_{OM}N\ln N+N^{2}_{OM})$, where $N$ is the number of photons in an OM, and the subsequent graph construction is quadratic in the number of OMs. As usually $N\gg N_{OM}$, the first term dominates, making it prohibitively costly for events that trigger a large portion of the array.
III Architectures
In our work, we use two different architectures for our classification problem; the first one uses kernel methods, while the second uses a Quantum Convolutional Neural Network (QCNN). In this section, we will describe these two algorithms.
III.1 Neural projected quantum kernels
Figure 2: Distribution of encoded variables. Each column represents a different energy range from 100 GeV to 1 PeV in four decades. Lighter dots corresponds to Track events, while darker dots are Cascade events. Top row shows $I_{1}-I_{0}$ as a function of $I_{2}-I_{1}$, while bottom row shows CoM as a function of these two variables. We display 300 data points for clarity. In our analysis, we exploit the separation shown in the bottom figures.
Neural quantum kernels were first introduced in Ref. [32] and later applied to a real-world classification problem in Ref. [5]. Kernel methods have several advantages over explicit methods; according to the representer theorem, kernel methods given the same encoding and dataset, achieve a lower or comparable training loss to explicit methods. In this work, we focus on neural projected quantum kernels (PQKs) where the inner product operations occur on a quantum computer.
The construction of neural PQKs begins with the training of an $n$-qubit quantum neural network (QNN). Since our goal is to implement the model on a real quantum computer, the experiments presented in the main text consider $n=2$. To build the 2-qubit QNN, we follow the iterative model proposed in Ref. [32]. This process starts by training a 1-qubit data re-uploading QNN [33], which consists of alternating layers of encoding and processing gates. The QNN architecture is defined as follows, acting on an input data point $\vec{x}\in\mathds{R}^{3}$ from the dataset
$\mathrm{QNN}_{\vec{\theta}}(\vec{x})=U(\vec{\theta}_{L})U(\vec{x})\dots U(\vec{\theta}_{1})U(\vec{x}),$ (2)
where $U(\phi,\theta,\omega)\equiv RZ(\omega)RY(\theta)RZ(\phi)$ represents a generic single-qubit rotation, and $RZ$ and $RY$ are single-qubit rotations about the $Z$ and $Y$ axes, respectively. The number of layers is denoted by $L$. The QNN is trained to optimize the $3L$ parameters $\vec{\theta}=\{\vec{\theta}_{1},\dots,\vec{\theta}_{L}\}$ by minimizing a cost function
$\vec{\theta}^{*}=\arg\min f_{\mathrm{cost}}(\vec{\theta}).$ (3)
We use the fidelity cost function, defined for $M$ training points as
$f_{\mathrm{cost}}(\vec{\theta})=\frac{1}{M}\sum_{i=1}^{M}\Bigl{(}1-|\langle\phi^{i}_{l}|\mathrm{QNN}_{\vec{\theta}}(\vec{x}_{i})|0\rangle|^{2}\Bigr{)}^{2},$ (4)
where $|\phi_{l}^{i}\rangle$ represents the correct label state for the data point $\vec{x}_{i}$ (chosen to be either $|0\rangle$ or $|1\rangle$).
Once the 1-qubit QNN is trained, we proceed to construct the 2-qubit QNN. This is done by extending the data re-uploading architecture described in Equation 2 to two qubits. The parameters for the single-qubit rotations on the first and second qubits are denoted as $\vec{\theta}^{(1)}=\{\vec{\theta}^{(1)}_{1},\dots,\vec{\theta}^{(1)}_{L}\}$ and $\vec{\theta}^{(2)}=\{\vec{\theta}^{(2)}_{1},\dots,\vec{\theta}^{(2)}_{L}\}$, respectively. Additionally, two-qubit gates are introduced at the end of each layer. These gates correspond to generic controlled rotations, where the control is on the second qubit
$CU(\vec{\varphi}_{i})=\mathds{1}\otimes|0\rangle\langle 0|+U(\vec{\varphi}_{i})\otimes|1\rangle\langle 1|.$ (5)
To implement the iterative training strategy, the 2-qubit QNN is initialized as follows: the parameters for the single-qubit gates of the first qubit are set to the optimal values obtained from the 1-qubit QNN training, i.e., $\vec{\theta}^{(1)}_{i}=\vec{\theta}^{*}_{i}$ for $i=1,\dots,L$, while the parameters for the two-qubit gates are initialized to zero, i.e., $\vec{\varphi}_{i}=\vec{0}$ for $i=1,\dots,L$. The parameters for the single-qubit gates of the second qubit, $\vec{\theta}_{i}^{(2)}$, can be initialized randomly. After this initialization, the optimization is performed over all parameters. Upon completion of the training, we obtain the 2-qubit quantum embedding $\mathrm{QNN}^{(2)}_{\vec{\theta}^{*},\vec{\varphi}^{*}}(\cdot)$.
Figure 3: Schematics of quantum architectures. (a) Neural projected quantum kernel construction. The embedding is constructed using the optimal parameters $\boldsymbol{\theta}$ obtained from training a QNN. The reduced density matrices of the first qubit are stored on a classical computer and used to construct the kernel entries $k_{ij}$, which are subsequently fed into a SVM for prediction. (b) Quantum convolutional neural network. Input data is amplitude-encoded and processed through convolutional layers consisting of parameterized unitaries. Pooling is performed via weighted measurements, effectively halving the number of qubits at each step. A final fully connected layer precedes measurement, enabling parameter optimization through a classical optimizer.
Finally, the 2-qubit neural PQK is constructed, as shown in  Figure 3 (a). The quantum embedding obtained from the trained QNN is used to define the kernel as the Hilbert-Schmidt inner product between the reduced density matrices of the first qubit
$k(\vec{x}_{i},\vec{x}_{j})=\mathrm{tr}\left(\rho_{(1)}(\vec{x}_{j})\rho_{(1)}(\vec{x}_{i})\right),$ (6)
where the reduced density matrix of the first qubit is given by:
$\rho_{(1)}(\vec{x}_{i})=\mathrm{tr}_{2}\left(\mathrm{QNN}_{\vec{\theta}^{*},\vec{\varphi}^{*}}^{(2)}(\vec{x}_{i})|00\rangle\langle 00|\mathrm{QNN}_{\vec{\theta}^{*},\vec{\varphi}^{*}}^{(2)}(\vec{x}_{i})^{\dagger}\right).$ (7)
This defines the kernel matrix $K$ with elements $k_{ij}=k(\vec{x}_{i},\vec{x}_{j})$. For further details on the construction and training methodology of neural quantum kernels, we refer the reader to the original work in Ref. [32]. One key aspect of this quantum kernel is that, since each kernel entry is computed using single-qubit reduced density matrices, the training phase can be less demanding in terms of quantum processing unit (QPU) usage. Instead of estimating all $M(M-1)/2$ pairwise fidelities on a quantum device—as is required in fidelity quantum kernels—only the $M$ reduced density matrices associated with the training points need to be obtained. Each of these $2\times 2$ matrices can be reconstructed from three single-qubit Pauli expectation values and stored compactly on a classical computer, where all pairwise Hilbert–Schmidt inner products can then be computed without further quantum processing.
Once the kernel matrix $K$ has been constructed, a classical support vector machine (SVM) is used to find the support vector coefficients $\vec{\alpha}$ which define the decision function
$f_{\vec{\alpha},X}(\vec{x})=\sum_{i=1}^{M}\alpha_{i}\;k(\vec{x},\vec{x}_{i}),$ (8)
where $X=\{\vec{x}_{i}\}_{i=1}^{M}$ denotes the training set. This function is then used to classify new data points by evaluating its sign: positive values correspond to one class and negative values to the other, in the binary classification setting.
Note that this construction also leads to substantial QPU savings in the prediction phase. In fidelity-based quantum kernels [34], evaluating the decision function for $M_{t}$ test points requires $M\times M_{t}$ fidelity estimates on a quantum processor. In contrast, PQKs only require adding the $M_{t}$ reduced density matrices corresponding to the test points to the previously stored $M$ reduced density matrices for the training points, after which the inner products can be computed classically.
However, given the current limitations of quantum processors—such as gate noise, decoherence, and restricted qubit connectivity—training the QNN and constructing the kernel entirely on real hardware remains out of reach. To nevertheless assess the feasibility of our method on actual quantum devices, we adopted a hybrid implementation strategy: the QNN was trained classically, while the reduced density matrices used for kernel construction were reconstructed experimentally on a quantum computer.
Hardware implementation — Experiments on real quantum hardware were carried out using IBM Strasbourg QPU, which features 127 qubits. To parallelize the acquisition of reduced density matrices, we selected 40 disjoint qubit pairs and measured their corresponding reduced density matrices simultaneously in each run. Figure 4 (c) shows the layout of the device, with the selected 40 qubit pairs highlighted. The pairs were chosen to be as physically distant as possible to reduce crosstalk and minimize hardware-induced correlations.
To reconstruct the reduced density matrices on the QPU, we measured the single-qubit Pauli expectation values $m_{i}=\langle\sigma_{i}^{(1)}\rangle$ for $i=X,Y,Z$, where the superscript 1 indicates that the measurement was performed on the first qubit of each pair. Each expectation value depends on the input data, i.e., $m_{i}=m_{i}(\vec{x})$, but we omit this dependence in the notation for simplicity. These expectation values are given by
$\langle\sigma_{i}^{(1)}\rangle=\langle 00|\mathrm{QNN}_{\vec{\phi}^{*},\vec{\varphi}^{*}}^{(2)}(\vec{x})^{\dagger}\;\sigma_{i}^{(1)}\;\mathrm{QNN}_{\vec{\phi}^{*},\vec{\varphi}^{*}}^{(2)}(\vec{x})|00\rangle,$ (9)
and were estimated using 1024 shots per observable. The reduced density matrix for the first qubit is then reconstructed as
$\rho_{(1)}(\vec{x})=\frac{1}{2}(\mathds{1}+m_{X}\;\sigma_{X}+m_{Y}\;\sigma_{Y}+m_{Z}\;\sigma_{Z}).$ (10)
Once the reduced density matrices for all training and test data points were generated, they were stored on a classical computer. During training, the Hilbert–Schmidt inner products between all training points were computed. For prediction, the Hilbert–Schmidt inner products between each test point and all training points were evaluated.
III.2 Quantum convolutional neural networks
We opted to implement a QCNN motivated by the demonstrated success of classical CNNs in classifying track and cascade events in IceCube data [28]. The underlying rationale is that, when a classical model performs well on a specific task, like event classification, the corresponding quantum model may also achieve comparable performance. Furthermore, it is well recognized that many quantum machine learning models suffer from trainability issues, such as the barren plateau problem, which severely hampers gradient-based optimization [35]. One advantage that QCNNs have over other variational quantum circuits is that they are barren-plateau-problem free [36], thus making them compelling candidates for problems involving high-dimensional input data. As illustrated in Figure 1 (c), our events can contain up to millions of photons, necessitating a model architecture –such as the QCNN– that is well-suited to handling high-dimensional data.
However, our initial implementations of the QCNN, which converted IceCube events into bit-strings using the data-encoding procedure proposed in an earlier work on quantum compression in neutrino astronomy [16], did not lead to effective training. We explored several alternative data mappings, including graph-based encodings and two-dimensional projections of IceCube event features, yet none led to significant separability. These early experiments highlighted the importance of introducing an appropriate inductive bias during our data encoding. Our results suggest that without physically motivated structure in the input representation, QCNN performance is limited given current available quantum hardware.
Following extensive experimentation with alternative encoding strategies, we adopted a moment of inertia based representation to map IceCube event data into our QCNN framework. This approach is grounded in strong physical intuition, as we outlined in earlier sections, and reduces the data feature space enough for us to easily simulate our QCNN — enabling rapid testing.
With the data representation established, the actual mapping of our moments of inertia into a quantum circuit used amplitude encoding. In this encoding scheme the data point $\vec{x}_{i}$ is first normalized, $\hat{x}$ to unity. Then given a set of n-qubits we can construct a Hilbert space of $2^{n}$ basis vectors, $|{\phi_{i}}\rangle$. We encode the data point $\vec{x}_{i}$ as a state $|\vec{x}_{i}\rangle$, given by
$|\vec{x}_{i}\rangle=\sum_{j}\hat{x}_{i}|\phi_{j}\rangle,$ (11)
i.e., $\langle\phi_{i}|\vec{x}_{i}\rangle=\hat{x}_{i}$. After the encoding step, the overall architecture adheres to the framework established by Cong et al. in their seminal QCNN paper [8] and is outlined in Figure 3. We apply a sequence of convolutional and pooling layers, repeated $d$ times, where $d$ denotes the QCNN depth. The convolutional layers use parameterized unitaries, $U(\vec{\theta})$ of the form:
$U(\vec{\theta})=e^{\sum^{N}_{i}(\theta_{i},\Lambda_{i})},$ (12)
where $\vec{\theta}_{i}$ is the set of trainable weights and $\Lambda_{i}$ are the $a\times a$ Gell-Mann matrices where $a=2^{w}$ and $w$ is the number of qubits involved in the unitary transformation. For the first convolutional layer, we apply unitaries to the set of all possible pairs of qubits, while later layers use multi-qubit unitaries. The pooling layers reduce the dimensionality of the system by measuring out half of the qubits in that layer; each measurement outcome informs a conditional arbitrary single qubit $V(\vec{\theta})$ where $V(\vec{\theta})\equiv V(\phi,\theta,\omega)\equiv RZ(\omega)RY(\theta)RZ(\phi)$. Once the desired depth $d$ is achieved, a fully connected layer of the same form of $U(\vec{\theta})$ is applied to the remaining qubits. At the output of the fully connected layer, we measure one of the remaining qubits to compute the predicted label.
This prediction is then used to evaluate our Mean-Squared Error (MSE) loss function which is optimized over by the classical optimizer Adam [37]. For each epoch in training, this classical optimization procedure informs the training weights how to change.
IV Results
In this section, we assess the performance of our quantum machine learning algorithms in classifying the light patterns produced by neutrino interactions in neutrino telescopes. Specifically, we aim to distinguish the elongated light patterns produced by muon neutrino interactions (tracks) from the more compact patterns resulting from other neutrino flavors (cascades). The dataset used for this study is described in Section II, and the quantum models are introduced in Section III. Each model’s parameters were optimized until convergence of the training loss.
Due to the current limitations of quantum hardware, encoding the full high-dimensional input data is infeasible. To address this, we employ physics-inspired feature engineering to reduce the data to three informative features per event. We explore two feature sets: (a) the three principal moments of inertia, $\vec{x}=(I_{2},I_{1},I_{0})$, and (b) two moment differences and the center-of-mass (CoM) traveled distance, $\vec{x}=(I_{2}-I_{1},I_{1}-I_{0},\texttt{CoM})$. These engineered representations, visualized in Figure 2, exhibit clear class separation, particularly in the 3D projections, thereby simplifying the classification task. Much of the model performance can thus be attributed to the discriminative power of this compact feature space, rather than to algorithmic complexity alone.
We conducted numerical experiments using the NPQK approach. A 2-qubit QNN with depth $L=10$ was used to generate the embedding. Each NPQK model was trained on 200 events and tested on 120, using the Adam optimizer with a learning rate of 0.05, a batch size of 8, and 20 training epochs. After training, the QNN served as a quantum feature map for kernel computation. The kernel matrix $K$ was evaluated under two settings: ideal simulation (purple curve) and execution on IBM’s 127-qubit Strasbourg quantum processor (blue curve). To enhance scalability, we used disjoint qubit pairs to parallelize the estimation of reduced density matrices, as described in section III.1. Each matrix was reconstructed from 1024 measurements per Pauli observable.
Figure 4: Classification results for tracks vs. cascades across different energy ranges using different feature sets. (a) Training and test accuracies using the three raw moments of inertia as input features, $\vec{x}=(I_{2},I_{1},I_{0})$. (b) Results using a feature set composed of two moment differences and the center-of-mass, $\vec{x}=(I_{2}-I_{1},I_{1}-I_{0},\texttt{CoM})$. For both feature sets, models are trained on 200 events and tested on 120. Metrics are reported as the mean and standard deviation across 5 random dataset splits. Results include: neural projected quantum kernels (NPQK) evaluated both in ideal simulation and on IBM’s Strasbourg QPU, and a quantum convolutional neural network (QCNN) evaluated under ideal simulation. (c) Visualization of the 40 qubit pairs we used, shown in blue, selected from the 127-qubit IBM Strasbourg quantum device. The pairs were chosen to maximize physical separation, enabling the simultaneous sampling of 40 reduced density matrices and thereby reducing the overall QPU runtime.
The QCNN was simulated with a width of 3 qubits and a depth of $L=5$. It was trained on 900 events, with 100 reserved for testing. Like the NPQK, optimization was performed using Adam, but with a cosine learning rate scheduler (initial rate 0.1, 100 decay steps, $\alpha=0.95$). Training and inference were carried out entirely in simulation, as the hybrid classical-quantum optimization loop remains impractical on current hardware.
Figure 4 summarizes classification results across four energy ranges for both the QCNN and NPQK models. The primary trend observed in the results is that both models exhibit improved performance with increasing energy. This behavior can be attributed to a physical effect: at lower energies, the geometric size of events becomes smaller, and when these event sizes approach the spatial resolution of the detector, as is the case at the lowest energies, distinctive morphological differences between tracks and cascades are largely lost. As a result, it becomes increasingly challenging for the models to distinguish between the two event categories as the event energy decreases.
A comparison between the NPQK and the QCNN shows that the NPQK consistently outperforms the QCNN. The QCNN exhibits difficulty fitting the data, as indicated by its persistently low standard deviation in training accuracy. This consistent lower performance suggests that the QCNN is not as apt at modeling the data, regardless of the training split, likely due to its limited expressivity. Although various model parameters, such as circuit width and depth, were tested for the QCNN, none led to improved test results.
While ideal simulation yields the highest accuracies overall, real hardware results are highly competitive. Notably, for the three highest energy bins, the test accuracy of the NPQK approach on real quantum hardware approaches 80%, closely matching ideal simulation. This is a remarkable result, especially considering the quantum circuit depth of 10 layers, which would typically be susceptible to noise and decoherence. The closeness of real and simulated performance underscores the robustness of the NPQK protocol, and in particular, validates the hybrid strategy where the quantum embedding is optimized classically, and only the quantum kernel is computed on the device. This separation allows the most noise-sensitive operations (training) to occur in a noise-free environment, while still leveraging the expressivity of quantum models during inference, thus demonstrating the practicality of current quantum resources for nontrivial classification tasks.
In the low-energy regime (below 10 TeV), model performance declines, with test accuracy reaching approximately 65%, despite a training accuracy of $\sim$75%. Compared to classical machine learning algorithms within this energy range, a similar accuracy is observed: a Boosted Decision Tree (BDT) achieves an accuracy of 67%, while a GNN attains 71% [24]. It is important to emphasize that the preprocessing and data cleaning procedures employed for these classical algorithms are considerably more extensive. For instance, they include requirements such as a minimum number of non-isolated hits within the signal region, an event vertex contained in a detector volume, starting/stopping positions of events near the detector, etc.; by contrast, our approach applies only a minimal cut on the number of hits (three or more). These results indicate that, at least for events containing limited information like those found at low energies ($\leq$ 1 TeV), the performance of quantum and classical approaches is comparable. The observed equivalence can largely be attributed to the dimensionality reduction technique adopted here, which allows our classification problem to be addressed using a feasible number of qubits while still maintaining much in the way of expressibility. These findings highlight the critical role of physics-informed feature construction in quantum machine learning. Given current constraints on qubit count and circuit depth, constructing low-dimensional yet informative input representations is key to achieving strong performance. The competitive results observed on real quantum hardware underscore the value of careful data preprocessing in enabling practical quantum classification.
V Conclusion
In this work, we present the first application of quantum machine learning to neutrino astronomy. We demonstrate that both Quantum Convolutional Neural Networks (QCNNs) and Neural Projected Quantum Kernels (NPQKs) can be successfully trained to classify neutrino events, specifically distinguishing track and cascade morphologies, with classification performance comparable to leading classical machine learning methods.
A significant success of our study lies in the formulation of a moment-of-inertia-based data encoding scheme. This physics-inspired representation enables quantum models to efficiently learn and establishes a template that can be readily extended to other datasets in neutrino astronomy.
Furthermore, we introduce a preprocessing strategy that dramatically reduces computational complexity. Unlike time quantile-based methods, which scale at least quadratically with the number of Optical Modules (OMs), our approach scales linearily with the number of OMs. This advantage is especially valuable for applications to large-scale neutrino observatories.
[…]
Although full-scale application remains constrained by the current state of quantum computers, our results underscore the promising future of hybrid quantum-classical approaches for machine learning tasks in neutrino astronomy. As quantum technologies continue to mature, we anticipate that quantum machine learning will play an increasingly valuable role in addressing complex scientific problems in neutrino astronomy and beyond.
Acknowledgments
We would like to thank Sean Chisholm for their help on the development of the earlier stages of the code used in this work. We would also like to thank Brad Baxley for their help making the cascade and track event display. PZ is supported by the David & Lucile Packard Foundation. CAA are supported by the Faculty of Arts and Sciences of Harvard University, the National Science Foundation (NSF), the NSF AI Institute for Artificial Intelligence and Fundamental Interactions, the Research Corporation for Science Advancement, and the David & Lucile Packard Foundation. CAA and PZ also acknowledge kind support and enriching environment of the Harvard Quantum Initiative; in particular, in the support of undergraduate researchers involved in the earlier stages of this project. PRG has received funding from the Basque Government through the “Plan complementario de comunicación cúantica” (EXP.2022/01341) (A/20220551). MS and PRG acknowledge support from HORIZON-CL4-2022-QUANTUM01-SGA project 101113946 OpenSuperQ-Plus100 of the EU Flagship on Quantum Technologies, the Spanish Ramón y Cajal Grant RYC-2020-030503-I, and the “Generación de Conocimiento” project Grant No. PID2021-125823NA-I00 funded by MICIU/AEI/10.13039/501100011033, by “ERDF Invest in your Future” and by FEDER EU. We also acknowledge support from the Basque Government through Grants No. IT1470-22, the Elkartek project KUBIT KK-2024/00105, and from the IKUR Strategy under the collaboration agreement between Ikerbasque Foundation and BCAM on behalf of the Department of Education of the Basque Government. This work has also been partially supported by the Ministry for Digital Transformation and the Civil Service of the Spanish Government through the QUANTUM ENIA project call – Quantum Spain project, and by the European Union through the Recovery, Transformation and Resilience Plan – NextGenerationEU within the framework of the Digital Spain 2026 Agenda. We acknowledge the use of IBM Quantum services for this work. The views expressed are those of the authors, and do not reflect the official policy or position of IBM or the IBM Quantum team.